OpenStudy (kirbykirby):

Moment generating functions: How can I show that \[Var(X)=\frac{d^2}{dt^2}ln M_X(t)\big |_{t=0}\]

OpenStudy (kirbykirby):

Recall: \(M_X(t)=E(e^{tx})=\int_{-\infty}^{\infty}e^{tx}f(x)dx\) \(E(X^n)=\frac{d^n}{dt^n}M_X(t)\big |_{t=0}\) \(Var(X)=E(X^2)-[E(X)]^2=E[(X-E(X))^2]\) ------------ I tried just applying the equation given but I don't know what to do with the log of this general integral? \[\frac{d^2}{dt^2}ln M_X(t)=\frac{d^2}{dt^2}ln \left( \int_{-\infty}^{\infty}e^{tx}f(x)dx) \right)\]

OpenStudy (kirbykirby):

Ah nvm the proof is way easier than I thought, and no integrals involved :)